OpenStudy (anonymous):

Solve the equation. If your answer is a fraction, use '/' to represent the fraction bar. For example, write 1/5 as 1/5 C- 2/12= 3/9 c=?

OpenStudy (kira_yamato):

C = 3/9 - 2/12 C = 1/3 - 1/6 = 1/6

OpenStudy (anonymous):

3/9 c = 3/9 * 1/6 = 1/18